OpenStudy (anonymous):

1/(1 - sinx) + 1/(1 + sinx) = 2sec^2x Prove the following identity, show all work.

OpenStudy (anonymous):

take LHS, and take LCM

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

uhh, i don't even know what the first person is saying. and um, thanks for the welcome. I've done \[sinx + 1 + sinx - 1\div(1-sinx)(1+sinx) = 2\div \cos^2x\] \[2sinx \div \sin^2x = 2 \div \cos^2x\] \[2sinx \div \cos^2x = 2 \div \cos^2x\] I dunno what I'm supposed to do for that last part? Did I mess up somewhere?

OpenStudy (anonymous):

\[LHS\] \[\frac{1}{1-sinx}+\frac{1}{1+sinx}\] taking LCM \[\frac{1+sinx+1-sinx}{(1-sinx)(1+sinx)}\] \[\frac{2}{\cos^{2}x}\] \[{2\sec^{2}x}\] \[=RHS\]